# Banyan: Improved Representation Learning with Explicit Structure

### v2

@@ 1 Introduction
To incorporate such inductive biases into models, the traditional information flow within neural networks needs altering. Instead of relying on implicit processing alone, models must learn representations for atomic components and an explicit computation graph that dictates how these components combine. Additionally, models must learn functions to govern information flow through this graph. Such approaches have demonstrated improved language modelling perplexity at cognitively plausible scales (Hu et al., 2021, 2022), better systematic generalisation (Sartran et al., 2022; Murty et al., 2023), and, especially relevant here, enhanced efficiency in acquiring semantics (Opper et al., 2023).
The Self-StrAE model (Opper et al., 2023) learns representations that explicitly model compositional semantics, and achieves promising performance while requiring minimal resources, both in terms of data and model size. It opened the door to exploring more compute-efficient solutions—particularly valuable for low-resource languages where scaling is often infeasible. However, while innovative, it still falls short compared to large-scale pre-trained transformers, even in languages outside standard pre-training corpora. Here, we introduce Banyan, a model which significantly outperforms Self-StrAE while achieving greater resource efficiency. Our approach involves modifying the structural optimisation process to induce an entangled graph that models global relations between nodes and employs a message passing mechanism using diagonal functions, reducing parameters while enhancing expressiveness.
@@ 2 Background and Related Work
Learning Structure: RvNNs often require structural input, which limits their flexibility since this structure may not always be available or easily obtainable. To address this, researchers have developed methods to induce structure within the model during recursive computation; using differentiable chart parsing (Drozdov et al., 2019, 2020; Hu et al., 2021, 2022), beam search (Ray Chowdhury & Caragea, 2023), continuous relaxation (Chowdhury & Caragea, 2021; Soulos et al., 2024), and reinforcement learning (Havrylov et al., 2019). However, these methods can struggle with memory issues and require careful tuning of hyperparameters. Here, we adopt a method from Opper et al. (2023) that uses representation similarity to determine how nodes should be merged during computation, which is both computationally efficient and surprisingly effective.
@@
Semantic Representation Learning through Structure: Transformer embeddings have become more successful than static word embeddings due to their ability to handle varying contextual influences. Unlike attention mechanisms in transformers, which route information based on token relationships, some approaches use explicit graphs or structures—such as dependency (Levy & Goldberg, 2014; Vashishth et al., 2019) or constituency (Opper et al., 2023) parses—to determine the focus of context windows. Opper et al. (2023) uses constituency parsers to learn sentence-level embeddings alongside word embeddings, introducing two models: StrAE, which relies on being given its structure, and Self-StrAE, which learns its own structure using representations. The latter model, Self-StrAE, serves as the foundation for Banyan and is described next.
@@ 3 Preliminary: Self-StrAE
Self-StrAE involves three main components that act over a sequence of tokens $\mathbf{w}=\langle w_{n}\rangle_{n=1}^{N}$: (a) an algorithm for merging tokens based on their similarity, (b) functions for composition and decomposition of embeddings, and (c) an objective that leverages both the induced structure and embeddings. While full details are available in Opper et al. (2023), we provide a brief overview to establish context for our model development (§ 4).
@@
To learn this model from data, Opper et al. (2023) derive two objectives. The first is straightforward cross-entropy over reconstructed tokens, for sentence $\mathbf{w}$ and prediction $\widehat{\mathbf{w}}$ as $\mathcal{L}_{\text{CE}}(\mathbf{w},\widehat{\mathbf{w}})=-\frac{1}{N}\sum_{n=1}^{N}w_{n}\cdot\log\widehat{w}_{n}.$ This however, places little constraint on the intermediate nodes in the hierarchical model. To address this, an alternate structural contrastive objective is formulated over a batch of sentences. As up and down trees are structurally identical (modulo edge reversal), it draws together an embedding and its dual on the other tree, while pushing away all other embeddings across the batch, using cosine similarity. Denoting pairwise similarity matrix $A\in\mathbb{R}^{M\times M}$ between up and down embeddings over $M$ nodes in the batch, the objective is: $\mathcal{L}_{\text{CO}}(\bar{e},\underline{e})=\frac{-1}{2M}\sum_{i=1}^{M}\log\left(\sigma_{\!\tau}(A_{i\bm{\cdot}})\,\sigma_{\!\tau}(A_{\bm{\cdot}i})\right)$ with $\sigma_{\!\tau}(\cdot)$ the tempered softmax over the unspecified ($\bm{\cdot}$) dimension.
@@ 5.1 Warmup: English language evaluation
Baselines: We compare against the Self-StrAE, GloVe embeddings (Pennington et al., 2014) and a RoBERTa (Liu et al., 2019) in the medium configuration from (Turc et al., 2019). Self-StrAE stands as the closest point of comparison to Banyan. Self-StrAE indicates the performance level of structured representation learning lies, as well as any improvements we are able to achieve. GloVe lets us compare to traditional static embeddings. This comparison probes whether our model is learning anything more than just simple bag of word features. To obtain sentence embeddings, we report results using both the simple average of the word embeddings and the average with stopwords removed following (Reimers & Gurevych, 2019). The latter is generally stronger. For RoBERTa, we report results using both the standard model, and again after enhancing RoBERTa through an extra round of contrastive SimCSE training (Gao et al., 2021), as a further STS baseline. Our pooling strategy is mean. To produce static embeddings from RoBERTa to use in lexical evaluation, we follow Bommasani et al. (2020) and average the contextualised representations of all occurrences of the word in the training set. The RoBERTa is intended as a stronger baseline. It has significantly more parameters than Banyan and is able to model meaning in context unlike GloVe.
@@
Hyperparameters and Pre-training Details: For all models we set the embedding size to 256. For Self-StrAE we use the configuration of (Opper et al., 2023) and set embeddings as square matrices (i.e., $K$=16 and $U$=16). For Banyan we set these values to $K$=128 and $U$=2, because the more independent channels we allowed the better the model seemed to perform. We refer the reader to the Appendix for ablations. We also note that because we can perform this reduction in channel size, the number of non-embedding parameters for Banyan drops to just 14, as these are directly proportional to $U$. We trained Self-StrAE and Banyan for 15 epochs (circa 15k steps and sufficient for convergence) using the Adam optimizer (Kingma & Ba, 2015), with a learning rate of 1e-3 for Banyan and 1e-4 for Self-StrAE  using a batch size of 512. To process the graphs we used DGL (Wang et al., 2020). The GloVe baseline was trained for 15 epochs with a learning rate of 1e-3, and a window size of 10. We used the official C++ implementation. RoBERTa medium was trained for 200,000 steps, (10% of which were used for warmup). We used a learning rate of 5e-5, and a linear schedule. Positional embeddings are relative key-query. The configuration for RoBERTa medium is 8 layers, 8 attention heads and 2048 dimensional feedforward layers. We used the Transformers library to implement and train the model (Wolf et al., 2020). For SimCSE training, we used the default parameters and the official implementation for unsupervised RoBERTa training from Gao et al. (2021). As our pre-training corpus we selected the WikiText-103 benchmark dataset (Merity et al., 2016). The RoBERTa and GloVe baselines are trained on the full corpus (103 million tokens), representing the upper-middle end of the level of data scale that might be available for a language. Whereas we trained Self-StrAE and Banyan on a uniform subsample of 10 million tokens, representing the lower end of how many tokens might be available, because these explicit structure models are supposed to be efficient learners.
@@ 5.4 Ablations
The simplest impacts stem from the introduction of the diagonalised composition and decomposition functions. These are sigmoided scalar values with which we multiply embeddings and act similarly to the fast weights of (Ba et al., 2016), decaying in the influence of embeddings further down in the tree on the root representation. This means that the embeddings produced by the model are restricted to conform to the compression order dictated by the structure, and we know from (Opper et al., 2023), that the more we can enforce this constraint the better our representations will end up. Secondly, such simple message passing functions bias the representation space towards informative separability. There has to be some signal from which to perform reconstruction, and all the pressure is now on the representations.
Switching the objective to cross entropy over the vocabulary rather than the contrastive objective used by (Opper et al., 2023) also yields significant benefits. This is likely because the contrastive loss is supposed to be beneficial because it enforces a pressure for representations to be uniformly distributed in space (Wang & Isola, 2020). However, our other modelling changes already push towards this quality. While it is a shame because the contrastive loss is conceptually elegant. It is known to have problems and eventually lead to shortcut solutions (Robinson et al., 2021). Therefore having a more robust objective grounded in data, like the cross entropy over the vocabulary, is actually quite nice.
@@ 6 Conclusion, Limitations and Future Work
Secondly, we have not yet fully exploited the potential of the inductive bias. Banyan still relies on greedy agglomerative clustering to induce structure. This is effective, but sub-optimal. Future work could learn the structure induction procedure. The type of structure models are exposed to impacts the quality of learnt semantic representations (Opper et al., 2023). So if how we induce structure improves, the model should learn better representations.
@@ Appendix A The k and u balance
The change to diagonal composition functions allows us to reduce the number of total parameters while maintaining performance. This is because the number of parameters is directly proportional to channel size u. We show ablations for this finding in Table 7. Our findings are similar to those of (Opper & Narayanaswamy, 2024) the smaller the channel size the better the model performs, although in our case we keep things stable between seeds whereas for them when they simplified they faced issues with extreme instability during training. This is thanks to the new message passing functions.

### v3

@@ 1 Introduction
To incorporate such inductive biases into models, the traditional information flow within neural networks needs altering. Instead of relying on implicit processing alone, models must learn representations for atomic components and an explicit computation graph that dictates how these components combine. Additionally, models must learn functions to govern information flow through this graph. Such approaches have demonstrated improved language modelling perplexity at cognitively plausible scales (Hu et al., 2021, 2022), better systematic generalisation (Sartran et al., 2022; Murty et al., 2023), and, especially relevant here, enhanced efficiency in acquiring semantics (Opper et al., 2023b).
The Self-StrAE model (Opper et al., 2023b) learns representations that explicitly model compositional semantics, and achieves promising performance while requiring minimal resources, both in terms of data and model size. It opened the door to exploring more compute-efficient solutions—particularly valuable for low-resource languages where scaling is often infeasible. However, while innovative, it still falls short compared to large-scale pre-trained transformers, even in languages outside standard pre-training corpora. Here, we introduce Banyan, a model which significantly outperforms Self-StrAE while achieving greater resource efficiency. Our approach involves modifying the structural optimisation process to induce an entangled graph that models global relations between nodes and employs a message passing mechanism using diagonal functions, reducing parameters while enhancing expressiveness.
@@ 2 Background and Related Work
Learning Structure: RvNNs often require structural input, which limits their flexibility since this structure may not always be available or easily obtainable. To address this, researchers have developed methods to induce structure within the model during recursive computation; using differentiable chart parsing (Drozdov et al., 2019, 2020; Hu et al., 2021, 2022), beam search (Ray Chowdhury & Caragea, 2023), continuous relaxation (Chowdhury & Caragea, 2021; Soulos et al., 2024), and reinforcement learning (Havrylov et al., 2019). However, these methods can struggle with memory issues and require careful tuning of hyperparameters. Here, we adopt a method from Opper et al. (2023b) that uses representation similarity to determine how nodes should be merged during computation, which is both computationally efficient and surprisingly effective.
@@
Semantic Representation Learning through Structure: Transformer embeddings have become more successful than static word embeddings due to their ability to handle varying contextual influences. Unlike attention mechanisms in transformers, which route information based on token relationships, some approaches use explicit graphs or structures—such as dependency (Levy & Goldberg, 2014; Vashishth et al., 2019) or constituency (Opper et al., 2023b) parses—to determine the focus of context windows. Opper et al. (2023b) uses constituency parsers to learn sentence-level embeddings alongside word embeddings, introducing two models: StrAE, which relies on being given its structure, and Self-StrAE, which learns its own structure using representations. The latter model, Self-StrAE, serves as the foundation for Banyan and is described next.
@@ 3 Preliminary: Self-StrAE
Self-StrAE involves three main components that act over a sequence of tokens $\mathbf{w}=\langle w_{n}\rangle_{n=1}^{N}$: (a) an algorithm for merging tokens based on their similarity, (b) functions for composition and decomposition of embeddings, and (c) an objective that leverages both the induced structure and embeddings. While full details are available in Opper et al. (2023b), we provide a brief overview to establish context for our model development (§ 4).
@@
To learn this model from data, Opper et al. (2023b) derive two objectives. The first is straightforward cross-entropy over reconstructed tokens, for sentence $\mathbf{w}$ and prediction $\widehat{\mathbf{w}}$ as $\mathcal{L}_{\text{CE}}(\mathbf{w},\widehat{\mathbf{w}})=-\frac{1}{N}\sum_{n=1}^{N}w_{n}\cdot\log\widehat{w}_{n}.$ This however, places little constraint on the intermediate nodes in the hierarchical model. To address this, an alternate structural contrastive objective is formulated over a batch of sentences. As up and down trees are structurally identical (modulo edge reversal), it draws together an embedding and its dual on the other tree, while pushing away all other embeddings across the batch, using cosine similarity. Denoting pairwise similarity matrix $A\in\mathbb{R}^{M\times M}$ between up and down embeddings over $M$ nodes in the batch, the objective is: $\mathcal{L}_{\text{CO}}(\bar{e},\underline{e})=\frac{-1}{2M}\sum_{i=1}^{M}\log\left(\sigma_{\!\tau}(A_{i\bm{\cdot}})\,\sigma_{\!\tau}(A_{\bm{\cdot}i})\right)$ with $\sigma_{\!\tau}(\cdot)$ the tempered softmax over the unspecified ($\bm{\cdot}$) dimension.
@@ 5.1 Warmup: English language evaluation
Baselines: We compare against the Self-StrAE, GloVe embeddings (Pennington et al., 2014) and a RoBERTa (Liu et al., 2019) in the medium configuration from (Turc et al., 2019; Opper et al., 2023a). Self-StrAE stands as the closest point of comparison to Banyan. Self-StrAE indicates the performance level of structured representation learning lies, as well as any improvements we are able to achieve. GloVe lets us compare to traditional static embeddings. This comparison probes whether our model is learning anything more than just simple bag of word features. To obtain sentence embeddings, we report results using both the simple average of the word embeddings and the average with stopwords removed following (Reimers & Gurevych, 2019). The latter is generally stronger. For RoBERTa, we report results using both the standard model, and again after enhancing RoBERTa through an extra round of contrastive SimCSE training (Gao et al., 2021), as a further STS baseline. Our pooling strategy is mean. To produce static embeddings from RoBERTa to use in lexical evaluation, we follow Bommasani et al. (2020) and average the contextualised representations of all occurrences of the word in the training set. The RoBERTa is intended as a stronger baseline. It has significantly more parameters than Banyan and is able to model meaning in context unlike GloVe.
@@
Hyperparameters and Pre-training Details: For all models we set the embedding size to 256. For Self-StrAE we use the configuration of (Opper et al., 2023b) and set embeddings as square matrices (i.e., $K$=16 and $U$=16). For Banyan we set these values to $K$=128 and $U$=2, because the more independent channels we allowed the better the model seemed to perform. We refer the reader to the Appendix for ablations. We also note that because we can perform this reduction in channel size, the number of non-embedding parameters for Banyan drops to just 14, as these are directly proportional to $U$. We trained Self-StrAE and Banyan for 15 epochs (circa 15k steps and sufficient for convergence) using the Adam optimizer (Kingma & Ba, 2015), with a learning rate of 1e-3 for Banyan and 1e-4 for Self-StrAE  using a batch size of 512. To process the graphs we used DGL (Wang et al., 2020). The GloVe baseline was trained for 15 epochs with a learning rate of 1e-3, and a window size of 10. We used the official C++ implementation. RoBERTa medium was trained for 200,000 steps, (10% of which were used for warmup). We used a learning rate of 5e-5, and a linear schedule. Positional embeddings are relative key-query. The configuration for RoBERTa medium is 8 layers, 8 attention heads and 2048 dimensional feedforward layers. We used the Transformers library to implement and train the model (Wolf et al., 2020). For SimCSE training, we used the default parameters and the official implementation for unsupervised RoBERTa training from Gao et al. (2021). As our pre-training corpus we selected the WikiText-103 benchmark dataset (Merity et al., 2016). The RoBERTa and GloVe baselines are trained on the full corpus (103 million tokens), representing the upper-middle end of the level of data scale that might be available for a language. Whereas we trained Self-StrAE and Banyan on a uniform subsample of 10 million tokens, representing the lower end of how many tokens might be available, because these explicit structure models are supposed to be efficient learners.
@@ 5.4 Ablations
The simplest impacts stem from the introduction of the diagonalised composition and decomposition functions. These are sigmoided scalar values with which we multiply embeddings and act similarly to the fast weights of (Ba et al., 2016), decaying in the influence of embeddings further down in the tree on the root representation. This means that the embeddings produced by the model are restricted to conform to the compression order dictated by the structure, and we know from (Opper et al., 2023b), that the more we can enforce this constraint the better our representations will end up. Secondly, such simple message passing functions bias the representation space towards informative separability. There has to be some signal from which to perform reconstruction, and all the pressure is now on the representations.
Switching the objective to cross entropy over the vocabulary rather than the contrastive objective used by (Opper et al., 2023b) also yields significant benefits. This is likely because the contrastive loss is supposed to be beneficial because it enforces a pressure for representations to be uniformly distributed in space (Wang & Isola, 2020). However, our other modelling changes already push towards this quality. While it is a shame because the contrastive loss is conceptually elegant. It is known to have problems and eventually lead to shortcut solutions (Robinson et al., 2021). Therefore having a more robust objective grounded in data, like the cross entropy over the vocabulary, is actually quite nice.
@@ 6 Conclusion, Limitations and Future Work
Secondly, we have not yet fully exploited the potential of the inductive bias. Banyan still relies on greedy agglomerative clustering to induce structure. This is effective, but sub-optimal. Future work could learn the structure induction procedure. The type of structure models are exposed to impacts the quality of learnt semantic representations (Opper et al., 2023b). So if how we induce structure improves, the model should learn better representations.
@@ Appendix A The k and u balance
The change to diagonal composition functions allows us to reduce the number of total parameters while maintaining performance. This is because the number of parameters is directly proportional to channel size u. We show ablations for this finding in Table 7. Our findings are similar to those of (Opper & Siddharth, 2024) the smaller the channel size the better the model performs, although in our case we keep things stable between seeds whereas for them when they simplified they faced issues with extreme instability during training. This is thanks to the new message passing functions.